Significance level of 0.05, test whether the slope of the regression line is negative.

The regression equation is, Y = 6.0 - 0.7x

predictor           coef               stdev          t-ratio

constant            6.0                2.558           2.32

X                      -0.7                0.086            -8.28

ANOVA

source           df                ss                    ms                F

regression      1               648.12                648.12           68.58

error               6                56.72                 9.45

total               7              704.84
Answer #1

ANSWER:

GIVEN THAT

  Significance level of 0.05

  The regression equation is, Y = 6.0 - 0.7 x

  • We need to test if the slope is negative.

       n = Total d+ 1 = 7 + 1 = 8

        n=8

  • Hence, Degrees of freedom for slope testing = n - 2 = 8 - 2 = 6
  • This is a left tailed test so critical t value at 0.05 significance level = -1.943
  • Obtained t value from output = -8.28 since -8.28 is less than the critical value, we reject Ho.
  • Therefore, we can conclude that the slope of the regression line is negative.